### v1

# Can dissipation induce a transition between many-body localized and thermal states?

Abstract
The many-body mobility edge (MBME) in energy, which separates thermal states from many-body localization (MBL) states, is a critical yet controversial concept in many-body systems. Here we examine the quasiperiodic $t_{1}-t_{2}$ model that features a mobility edge. With the addition of nearest-neighbor interactions, we demonstrate the potential existence of a MBME. Then we investigate the impact of a type of bond dissipation on the many-body system by calculating the steady-state density matrix and analyzing the transport behavior, and demonstrate that dissipation can cause the system to predominantly occupy either the thermal region or the MBL region, irrespective of the initial state. Finally, we discuss the effects of increasing system size. Our results indicate that dissipation can induce transitions between thermal and MBL states, providing a new approach for experimentally determining the existence of the MBME.
I Introduction
Closed quantum many-body systems with disorder or quasiperiodic potentials can exhibit localization review1 ; review2 , wherein the system cannot act as its own heat bath and thus does not reach thermodynamic equilibrium. Many-body localization (MBL) exhibits several intriguing properties, such as the absence of conductivity (even at finite temperatures) review2 ; Basko2006 ; Gornyi2005 , specific spectral properties Huse2007 ; Huse2010 ; Serbyn2016 , entanglement entropy that follows an area law review1 ; review2 , and the slow logarithmic growth of entanglement entropy Chiara2006 ; Prosen2008 ; Moore2012 ; Serbyn2013 . In addition to its fundamental theoretical importance, MBL has potential significant applications in quantum information Altman2015 ; Nayak2014 , time crystals NormanYao , and other areas, which has garnered widespread theoretical and experimental attention. For example, MBL phenomena have been observed in platforms such as cold-atom systems IBloch2015 ; IBloch2016 ; Islam2015 , circuits with superconducting qubits Roushan2017 ; HHWang2021 and trapped-ion JSmith2016 systems. However, even in these highly controlled implementations, MBL systems are influenced by at least a slight coupling to the environment, such as inelastic scattering from lasers, which may ultimately disrupt localization. Furthermore, considering the observation of MBL in traditional solid-state experiments Lake2207 and the applications of MBL phenomena, the impact of dissipation on MBL becomes even more unavoidable. Therefore, the impact of dissipation on MBL has attracted extensive research attention JSmith2016 ; DAHuse2014 ; Johri2015 ; DAHuse2015 ; Hyatt2017 ; Fischer2016 ; Levi2016 ; Medvedyeva ; Everest2017 ; Knap2017 ; LNWu2019 ; JRen2020 ; Gopalakrishnan ; Roeck2017 ; Luitz2017 ; Richter2024 ; Denisov2018 ; Bloch2017 ; Bloch2019 . When coupled to a thermal bath, dissipation can lead to infinite heating in the long-time limit, making MBL typically considered unstable DAHuse2014 ; Johri2015 ; DAHuse2015 ; Hyatt2017 ; Fischer2016 ; Levi2016 ; Medvedyeva ; Everest2017 ; Knap2017 ; LNWu2019 ; JRen2020 ; Gopalakrishnan ; Roeck2017 ; Luitz2017 , as observed in cold-atom experiments Bloch2017 ; Bloch2019 . Although dissipation can disrupt MBL, transport in the dissipation-induced delocalized phase is different from that in a typical thermalized phase Knap2017 . Additionally, once the dissipation is removed, the system returns to a many-body localized state. Can dissipation drive a transition from a MBL state to a thermal state? Conversely, can a certain type of dissipation drive a transition from a thermal state to a MBL state?
The concept of the mobility edge (ME) is a central idea in the field of localization physics. Similar to how MBL can be seen as the counterpart of Anderson localization in many-body systems, a natural question arises: can the concept of the ME be extended to many-body systems? In other words, is there a critical energy that separates thermal states from many-body localized states, known as the many-body mobility edge (MBME)? Numerous numerical results and some experiments suggest the existence of the MBME MBME0 ; MBME1 ; MBME2 ; MBME3 ; MBME4 ; MBME5 ; MBME6 ; MBME7 ; MBME8 . However, due to computational size limitations, the existence of a definitive MBME remains uncertain. For instance, Ref. noME argues that local fluctuations in the system with a putative MBME can act as mobile bubbles, inducing global delocalization, and hence, a MBME cannot exist. If a MBME exists, could the introduction of dissipation lead to some interesting results? Would such results provide insights into determining the MBME?
In this work, we demonstrate that by introducing a special type of controllable dissipation into a system, which, from the perspective of computable sizes, has the MBME, we can drive the system into a steady state, which predominantly consists of either thermal states or many-body localized states, independent of the initial state. Therefore, within the limits of computable finite sizes, we can assert that dissipation can induce transitions between thermal states and MBL states. At the same time, our results also provide a feasible approach for determining the existence of the MBME in simulated systems such as cold atoms.
II Model and Results
We consider the quasiperiodic $t_{1}-t_{2}$ model t1t2 with nearest-neighbor (NN) interactions, given by
$H$ $=$ $-\sum_{j}\left(t_{1}c_{j+1}^{{\dagger}}c_{j}+t_{2}c_{j+2}^{{\dagger}}c_{j}+\text{H.c.}\right)+U\sum_{j}n_{j}n_{j+1}$ (1)
$+2\lambda\sum_{j}\cos\left(2\pi\omega j+\phi\right)n_{j},$
where $t_{1}$ and $t_{2}$ are the NN and next-NN hopping amplitudes, respectively, $U$ is the NN interaction strength, $\omega$ is an irrational number, $\lambda$ is the quasiperiodic potential strength, and $\phi$ is a phase offset. Without loss of generality, we take $t_{1}=1$, $t_{2}=0.2$, and $\omega=\left(\sqrt{5}-1\right)/2$. Unless otherwise stated, we will use open boundary conditions in subsequent computations.
In the non-interacting limit ($U\rightarrow 0$), this model displays a single-particle ME, as shown in Fig. 1(a), where we present the inverse participation ratio (IPR) of each eigenstate, which, for an arbitrary $m$-th eigenstate $|\psi_{m}\rangle=\sum_{j}^{L}\psi_{m,j}c_{j}^{\dagger}|\varnothing\rangle$ with $L$ being the system size, is defined as $IPR=\sum_{j=1}^{L}\left|\psi_{m,j}\right|^{4}$. It is known that for extended and localized states, the IPR tends to 00 or a finite nonzero value, respectively. From Fig. 1(a), we see that the extended and localized states are divided by one ME. In general, the $t_{1}-t_{2}$ model does not have an analytical expression for the ME. However, when $t_{1}/t_{2}$ is sufficiently large, the NN and next-NN hopping terms can be approximated by an exponential hopping with strength $t=t_{1}e^{p}$ with $p=\ln(t_{1}/t_{2})$ Biddle1 ; Biddle2 . The latter has an exact expression for the ME, and based on this, we can obtain an approximate expression for the ME of the $t_{1}-t_{2}$ model in this case:
$E_{c}=\frac{-\lambda(t_{1}^{2}+t_{2}^{2})+t_{1}^{3}}{t_{1}t_{2}},$ (2)
as shown by the dashed line in Fig. 1(a).
When the NN interaction is included, the single-particle ME may lead to the emergence of a MBME. To characterize the localized and thermal properties of this system, we consider the ratio of adjacent energy gaps Huse2007 ; Huse2010 ; Serbyn2016 as $r_{i}=\frac{\min\left(\delta_{i+1},\delta_{i}\right)}{\max\left(\delta_{i+1},\delta_{i}\right)}$, where $\delta_{i}=E_{i}-E_{i+1}$ represents the energy spacing with the eigenvalues $E_{i}$ listed in ascending order. For the system in the thermal region, its level statistics follow the Wigner-Dyson distribution, and the average $\left\langle r\right\rangle$ converges to $0.53$. In the MBL region, the level statistics are Poisson, with $\left\langle r\right\rangle\approx 0.39$. We rescale the many-body spectrum as $\epsilon_{i}=\left(E_{i}-E_{g}\right)/\left(E_{\max}-E_{g}\right)$, where $E_{g}\left(E_{\max}\right)$ is the eigenenergy of the ground (highest excited) state. Based on this, we divide the eigenvalues into $10$ different energy windows and average samples and all gaps in each window to obtain an $\left\langle r\right\rangle$, as shown in Fig. 1(b). We observe that as the strength of the quasiperiodic potential $\lambda$ increases, the states in the uppermost window become localized first, similar to the single-particle case. In the inset of Fig. 1(b), we fix $\lambda=0.8$. The red and blue lines correspond to the $\left\langle r\right\rangle$ of the lowest energy window and the highest energy window, respectively, which converge to $0.53$ and $0.39$. This indicates that the states in the highest window become localized while the states in the lowest window remain thermalized, suggesting the existence of a MBME. In the following discussion, we fix $\lambda=0.8$ and one-third filling.
Figure 1: (a) IPR of various eigenstates plotted against their corresponding eigenvalues $E$ and the strength of the quasiperiodic potential of the $t_{1}-t_{2}$ model. Black dashed line indicates the approximate mobility edge obtained from Eq. (2). (b) Variation of $\left\langle r\right\rangle$ with energy density ($\epsilon$) and quasiperiodic potential strength $\lambda$. Here $t_{1}=1$ and $t_{2}=0.2$, $U=0.9$, the system size $L=18$, the number of particles $N=6$, and the number of samples is $20$, with each sample corresponding to an initial phase $\phi$. Inset: Variation of $\left\langle r\right\rangle$ with system size at one-third filling for the highest (blue) and lowest (red) energy window.
Next, we introduce the dissipation affecting a pair of neighboring sites $j$ and $j+1$, described by BD1 ; BD2 ; BD3 ; BD4 ; BD5 ; BD6 ; BD7 ; BD8 ; BD9
$O_{j}=(c_{j}^{{\dagger}}+ac_{j+1}^{{\dagger}})(c_{j}-ac_{j+1}),$ (3)
where $a=\pm 1$, and $j=1,2,\ldots,L-1$. This form of dissipation can be implemented using cold atoms in optical superlattices BD1 ; BD2 ; BD3 ; BD4 or through arrays of superconducting microwave resonators BD6 . This bond dissipation maintains the particle number but modifies the relative phase between adjacent lattice sites. It synchronizes them from an in-phase (out-of-phase) mode to an out-of-phase (in-phase) mode when $a$ is set to $-1$ ($1$).
The dissipative dynamics of density matrix $\rho$ is represented by the Lindblad master eqaution GLindblad ; HPBreuer
$\frac{\mathrm{d}\rho(t)}{\mathrm{d}{}t}=\mathcal{L}\left[\rho\left(t\right)\right]$ $=$ $-i\left[H,\rho\left(t\right)\right]$ (4)
$+\Gamma\sum_{j}\left(O_{j}\rho O_{j}^{{\dagger}}-\frac{1}{2}\left\{O_{j}^{{\dagger}}O_{j},\rho\right\}\right),$
Where $\mathcal{L}$ is the Lindbladian superoperator, and the strength of jump operators is set to $\Gamma$, which is independent of the lattice points. Our results are essentially independent of the value of $\Gamma$, and without loss of generality, we set $\Gamma=1$. Here, $\mathcal{L}$ is time-independent, so we can express $\rho(t)=e^{\mathcal{L}t}\rho(0)$. The steady state is defined as $\rho_{s}=\rho(t\rightarrow\infty)$, which corresponds to the eigenstate of the Lindbladian $\mathcal{L}$ with the zero eigenvalue, i.e., $\mathcal{L}[\rho_{s}]=0$. We analyze the properties of the steady-state density matrix $\rho_{s}$ in the eigenbasis of the many-body Hamiltonian $H$, that is, $\rho_{nm}=\langle\psi_{n}|\rho_{s}|\psi_{m}\rangle$, where $|\psi_{n}\rangle$ and $|\psi_{m}\rangle$ denote the eigenstates of $H$. Fig. 2 illustrates that the system’s steady state primarily occupies the low-energy thermal region when $a=1$ [Fig. 2(a)], whereas it predominantly occupies the high-energy MBL region when $a=-1$ [Fig. 2(b)]. This means that by adjusting dissipation, we can control whether the steady state of this system predominantly resides in the thermalization region or in the MBL region.
Figure 2: Absolute value of steady-state density matrix elements in the Hamiltonian eigenstate basis with (a) $a=1$ and (b) $a=-1$. Here we take $L=9$, $N=3$, $U=0.9$, $\lambda=0.8$, and $\Gamma=1$.
When dissipation is present, we note that even if the steady state of the system resides in a localized region under the basis of the Hamiltonian, the system’s dynamics are not necessarily localized. For instance, dissipation mechanisms like dephasing can disrupt coherence and consequently destroy the system’s localization properties. The bond dissipation described by Eq. (3) in our study expands to include terms related to dephasing dissipation, which also undermine localization. However, in the case of dephasing dissipation disrupting localization, once dissipation is removed, coherence and localization can be restored, completely eliminating the effects caused by dephasing. In contrast, the bond dissipation we discuss selectively stabilizes the system in specific regions through state selection. Upon removing dissipation, the effects it induces do not vanish. This can be observed through the evolution of the density matrix after dissipation is removed. After abruptly removing the dissipation, we observe the evolution of the density matrix: $\rho(t)=\sum_{mn}e^{i(E_{m}-E_{n})t}\rho_{mn}|\psi_{m}\rangle\langle\psi_{n}|$, where $E_{m}$ and $E_{n}$ are the eigenvalues corresponding to the many-body eigenstates $|\psi_{m}\rangle$ and $|\psi_{n}\rangle$. We observe that the diagonal elements of $\rho$ (where $E_{m}=E_{n}$) do not change with time, while the off-diagonal elements fluctuate over time. However, with longer observation times, we need to examine the average effects of the dynamics over this interval, leading to the vanishing of the effects caused by the off-diagonal elements. Therefore, when the steady state primarily occupies the localized (thermalization) region, removing dissipation reveals behavior characteristic of the localized (thermalization) region. Based on this observation, we can conclude that dissipation serves as an intermediate process capable of inducing a transition between thermalization and MBL states.
Experimental platforms such as ultracold atoms transportExp1 ; transportExp2 ; transportExp3 ; transportExp4 ; transportExp5 can probe the transport properties of a system. Next, we will examine the transport properties of the system after it reaches a steady state and the dissipation is removed. By using these properties to distinguish between thermalization and MBL behaviors, we can experimentally verify that dissipation can induce transitions between thermal and localized states. We assume the system reaches a steady state at the time $t_{0}$, after which dissipation is removed. At this point, the system returns to a non-equilibrium state, and we then study its response to a probed electric field. The change in current $\delta I$, for a weak probed electric field, can be derived from linear response theory.
[…]
where $\sigma(t,t^{{}^{\prime}})$ is the non-equilibrium conductivity, which depends only on the time difference $t-t^{{}^{\prime}}$ for an equilibrium state transport1 . Here we consider a delta probed field $E(t)=E\delta(t-t_{0})$ such that $\langle\delta I(t)\rangle=LE\sigma(t,t_{0})$. For convenience, we set $e=\hbar=E=1$. So the conductivity in a finite size system can be written as transport2
$\sigma(t,t^{{}^{\prime}})=\frac{i}{N}\theta(t-t^{{}^{\prime}})Tr[\rho(t_{0})[\hat{I}(t),\hat{B}(t^{{}^{\prime}})]],$ (6)
where the position operator $B=\sum_{j}jc_{j}^{\dagger}c_{j}$ and the current operator $I=\frac{dB}{dt}=-i[B,H]$. The current change in the thermal region is expected to show large fluctuations because it is significantly affected by the electric field, whereas in the MBL state, the current change should exhibit smaller fluctuations due to minimal influence from the electric field [see the appendix A]. In Fig. 3, the blue line represents the $a=1$ case. As discussed earlier, in this scenario, the steady state primarily occupies the thermalization region near the ground state. After removing dissipation, we observe that the current change consistently exhibits larger oscillations. In contrast, when $a=-1$ (the orange line in Fig. 3), the system’s steady state mainly occupies the many-body localized region near the highest state. After removing dissipation, we see that the current change oscillates with a smaller amplitude and evolves to near zero at a faster rate, exhibiting relatively localized properties.
Figure 3: The conductance of the steady-state density matrix after removing dissipation varies with time. The red line corresponds to the case of $a=-1$, where the steady state mainly occupies the localized region near the highest excited state. In contrast, the blue line corresponds to $a=1$, where the steady state mainly occupies the thermalized region near the ground state. Other parameters are the same as in Fig. 2.
III Discussion on the effects of system size
As discussed earlier, in small systems, there exists a MBME, and dissipation can induce transitions between thermal and MBL states. However, when the system size increases, the existence of a MBME becomes a controversial issue. If there is no MBME, then the dissipation-induced transitions we discussed here would not exist. We use the density matrix renormalization group (DMRG) method to study the thermalization and localization properties of the ground state and the highest excited state in large systems. To characterize the localization properties of these two states, we consider the density distribution of single-particle excitation, defined as the difference in distribution after introducing an additional particle into the many-body system, i.e., $\delta n_{j}=\rho_{N+1}\left(j\right)-\rho_{N}\left(j\right)$, where $\rho_{N}\left(j\right)=\langle\psi^{N}_{g(e)}|n_{j}|\psi^{N}_{g(e)}\rangle$ and $|\psi^{N}_{g(e)}\rangle$ is the ground (highest excited) state of the system with $N$ particles. We add a particle to a system with size $L=144$ and particle number $N=32$. Fig. 4 shows the distribution of $\delta n$ at each lattice site. We see that for the ground state (the orange dots), the single particle excitation is distributed quite uniformly across different lattice sites, indicating an extended property. In contrast, for the highest excited state (the blue dots), $\delta n_{j}$ is primarily concentrated on a few lattice sites, showing localized behavior. These results correspond to those in Fig. 1, indicating a significant difference in localization properties between the ground state and the highest excited state in larger systems. Similarly, we can also define the IPR of single-particle excitation (SIPR) as
$SIPR=\frac{\sum_{j}\delta n_{j}^{2}}{\sum_{j}|\delta n_{j}|}.$ (7)
When the distribution of single-particle excitation is extended or localized, the corresponding SIPR respectively tends to 00 or a finite nonzero value. The SIPR corresponding to the ground state and the highest excited state are $0.023$ (close to 00) and $0.301$, respectively [Fig. 4], further indicating that they are extended and localized, respectively.
Figure 4: The distribution of $\delta n$ at each lattice site after adding a particle to a system with $L=144$ and $N=32$ for the ground state (orange dots) and the highest excited state (blue dots), with their corresponding SIPR values being $0.023$ and $0.301$, respectively.
Our results suggest that the MBME might exist in large systems. In this case, the transition between thermal and MBL states induced by dissipation should also be present in large systems. This provides a potential method for experimentally determining whether the MBME exists. Although the MBME has been observed in some systems, such as superconducting qubit circuits HHWang2021 , the system sizes used in experiments are still not large enough. Cold atom systems can simulate larger systems, but measuring the MBME in the system presents significant challenges BlochME . Based on our results, by adjusting dissipation, we can selectively place the system in either the thermal region or the MBL region, facilitating measurements. This makes it possible to detect the MBME in cold atom experiments.
IV Conclusion
We have investigated the impact of a type of bond dissipation on the quasiperiodic $t_{1}-t_{2}$ model with NN interactions, which possesses a mobility edge in the absence of interactions. By calculating the level spacing distribution, we observed the potential existence of a MBME. We further analyzed the distribution of the steady-state density matrix and revealed that dissipation can drive the many-body system into specific states primarily located in either the thermal region or the MBL region, regardless of the initial states. Thus, dissipation can be used to induce transitions between thermal and MBL states, allowing for the control of particle transport behaviors, as demonstrated by our analysis of the system’s transport properties after removing dissipation. Finally, we used DMRG to study single-particle excitations in the ground state and highest excited state for large system sizes, demonstrating their different localization behaviors. Our findings suggest that dissipation can trigger transitions between thermal and MBL states, offering a novel method for experimentally identifying the presence of the MBME.
Acknowledgements
This work is supported by National Key R&D Program of China under Grant No.2022YFA1405800, the National Natural Science Foundation of China (Grant No.12104205), the Key-Area Research and Development Program of Guangdong Province (Grant No. 2018B030326001), Guangdong Provincial Key Laboratory (Grant No.2019B121203002).
Appendix A Change in current: An example using the AA model
We take the interacting Aubry-André (AA) model (i.e., with next-nearest-neighbor hopping strength $t_{2}=0$ in the Hamiltonian of Eq. (1)) as an example to illustrate the different current changes between thermal and many-body localized states after introducing an electric field. We still fix one-third filling. We consider only the ground state and select two extreme values of the quasiperiodic potential: $\lambda=0.1$, representing the thermal state property, and $\lambda=10$, representing the MBL state property.
As shown in Fig. 5, the current change in the thermal state has a large amplitude and a long oscillation period, similar to the case of $a=1$ shown in Fig. 3 of the main text, while the MBL state exhibits a small amplitude around 00 and fast oscillations, similar to the case of $a=-1$ shown in Fig. 3 of the main text. This can be understood as the current in the thermal state being significantly affected by the electric field, whereas the MBL state is minimally influenced by the electric field.
Figure 5: The current change $\delta I$ of the ground state for the many-body AA model for different quasiperiodic potential strengths: blue line for $\lambda=0.1$ and orange line for $\lambda=10$. Other parameters are $L=9$, $N=3$, $t_{1}=1$, $t_{2}=0$, and $U=1$.